### v3

# SAT: Dynamic Spatial Aptitude Training for Multimodal Language Models

Abstract
equal advising, joint second author
Reasoning about motion and space is a fundamental cognitive capability that is required by multiple real-world applications. While many studies highlight that large multimodal language models (MLMs) struggle to reason about space, they only focus on static spatial relationships, and not dynamic awareness of motion and space—i.e., reasoning about the effect of egocentric and object motions on spatial relationships. Manually annotating such object and camera movements is expensive. Hence, we introduce SAT, a simulated spatial aptitude training dataset utilizing 3D simulators, comprising both static and dynamic spatial reasoning across 175,000 question-answer (QA) pairs and 20,000 scenes. Complementing this, we also construct a small (150 image-QAs) yet challenging dynamic spatial test set using real-world images. Leveraging our SAT datasets and 6 existing static spatial benchmarks, we systematically investigate what improves both static and dynamic spatial awareness. Our results reveal that simulations are surprisingly effective at imparting spatial aptitude to MLMs that translate to real images. We show that perfect annotations in simulation are more effective than existing approaches of pseudo-annotating real images. For instance, SAT training improves a LLaVA-13B model by an average $11\%$ and a LLaVA-Video-7B model by an average $8\%$ on multiple spatial benchmarks, including our real-image dynamic test set and spatial reasoning on long videos—even outperforming some large proprietary models. While reasoning over static relationships improves with synthetic training data, there is still considerable room for improvement for dynamic reasoning questions.
Figure 1: We propose Spatial Aptitude Training (SAT), which generates simulated spatial training data that allows us to go beyond simple static relationships in existing datasets. Inspired by cognitive science, we add more challenging dynamic questions that require reasoning about ego and object motion. SAT improves performance on both existing spatial benchmarks and dynamic reasoning on our new real-image benchmark.
1 Introduction
Reasoning about space and motion is a fundamental cognitive capability that allows humans and animals to survive and operate in the real world (gardner2011frames; vasilyeva2012development; blades1994development). While simple visual relationships (such as left-right) may be trivial to recognize for humans, more dynamic reasoning involving motion such as taking perspectives and reasoning about the effect of ego-motion or object movement requires more cognitive support (gardner2011frames; tversky2019mind). Hence, it is not surprising that such “dynamic” spatial reasoning capabilities in children correlate with their aptitude in geometry, physics, and even linguistic reasoning (taylor1992spatial; mallot2024geometry).
Despite the promise of multimodal language models (MLMs) as general-purpose intelligent agents (achiam2023gpt; zhu2023minigpt; liu2024visual), recent studies reveal that these models still struggle with spatial reasoning (chen2024spatialvlm; cheng2024spatialrgptgroundedspatialreasoning; tong2024cambrian1fullyopenvisioncentric; fu2024blink). Specifically, MLMs often fail to predict the relative positions of objects in static images—a limitation attributed to the scarcity of spatial relationship annotations in their training data. To address this, recent methods (cheng2024spatialrgptgroundedspatialreasoning; chen2024spatialvlm) introduce pseudo-annotations to encode static object relationships in real images (cheng2024spatialrgptgroundedspatialreasoning), which require extensive engineering and refining to scale. Furthermore, many real-world applications require reasoning that extends beyond static object positions. For example, smart glasses and embodied AI applications need to reason about dynamic scenes, where reasoning about the effect of object and camera movements are essential (duan2024manipulate; yuan2024robopoint; liu2024coarse). Generating pseudo-annotations using existing methods for such dynamic settings is non-trivial due to the complexity of predicting action causality on real images.
We explore the possibility of using simulation data as a simple solution to impart both static and dynamic spatial reasoning into MLMs. We propose Spatial Aptitude Training (SAT), an approach to generate spatial question-answer (QA) data without any human supervision. Manually annotating 3D movements is expensive; hence, SAT leverages 22K ProcTHOR (deitke2022️) scenes composed of 1K assets to generate 175K QA pairs. With perfect 3D information and control of the assets, SAT goes beyond static object relationships to questions that require reasoning about egocentric and object movements. Since our data is generated procedurally by composing assets, it can be scaled up arbitrarily. Hence, SAT is programmatically controllable, unlike existing 3D (brazil2023omni3d; azuma2022scanqa) or spatial datasets (cheng2024spatialrgptgroundedspatialreasoning), which are also not compositional.
With SAT, we analyze what types of training data improve spatial reasoning. We focus on two types of spatial reasoning data. First, spatial-QA about object relations in static scenes, shown in Fig. 1 (under “Existing Spatial Bench”), to impart reasoning about the relative locations of objects in the scene (e.g., where is object X with respect to object Y?). Next, we evaluate the effect of dynamic spatial tasks, Fig. 1 (shown under “Our SAT Dynamic Bench”). This includes questions about egocentric movement, object movement, allocentric perspective, goal aiming, and action consequences. These dynamic spatial reasoning skills are well-studied in human cognitive development: children understand the consequences of self-motion (e.g., the moving room test (anderson2013role)), track how entities move in a scene (anderson2013role), and can adopt others’ viewpoints when planning actions (brucato2023measuring; vasilyeva2012development; blades1994development).
We use two widely adopted open-source MLMs, an image-based LLaVA-1.5-13B and a video-based LLaVA-Video-7B, as our base models for evaluations. To test static spatial reasoning, we use four contemporary real image benchmarks: CV-Bench (tong2024cambrian1fullyopenvisioncentric), BLINK (fu2024blink), Visual Spatial Relations (VSR) dataset (liu2023visualspatialreasoning), and GQA-Spatial (hudson2019gqanewdatasetrealworld). In addition, we test the effect of SAT on spatial reasoning on videos using the recent VSI-Bench (yang2024thinking) benchmark and MME-RealWorld (zhang2024mme), which include out-of-domain outdoor videos in realistic self-driving scenarios. Since no dataset exists for dynamic spatial reasoning on images (predicting action causality on images), we construct a difficult SAT test set on real images.
Challenging conventional wisdom about sim-to-real transfer in model training, we find simulations to be surprisingly effective at imparting spatial intelligence to MLMs. Our simulated SAT training improves the baseline LLaVA-13B model by $11\%$ and LLaVA-video-7B by $8\%$ (absolute) on average on a wide range of spatial benchmarks, including our challenging dynamic test set—even outperforming some larger closed-source and spatially-tuned models (achiam2023gpt; team2023gemini). Interestingly, SAT also improves spatial reasoning on long videos (yang2024thinking)—especially in route planning by $4\%$, showing promise for aiding embodied applications. However, while simple static relationships in existing datasets are easy to improve, our best model performs at $62\%$ on our SAT dynamic questions (random chance is $50\%$)—hence, leaving considerable room for improvement.
Our work shows promise that training with data generated using embodied movements and interactions in simulators can indeed help instill spatial intelligence in MLMs.
Table 1: Comparison of existing datasets to ours. Our training dataset is synthetic and interactive, allowing us to generate spatial QA and extend to dynamic (e.g., object movement) reasoning data for free. It can also be easily extended to generate new scenes and tasks.
SAT
2D Vision-Language
GQA, VG, Obj365
3D Vision-Langauge
Omni3D, ScanQA
Spatial Rel
VSR, 2.5VRD
Spatial QA
CVBench, BLINK,
Sp VLM, Sp RGPT
2D Annotations
3D Annotations (2.5D)
Static Spatial QA
Dynamic Spatial QA
New Scene/Task Gen
Object Interaction
2 Related work
Our work draws inspiration from fundamental schools of thought in neuroscience that suggest spatial intelligence is a core foundation for most cognitive abilities (tversky2009thinking; mallot2024geometry; duan2022survey).
Spatial understanding benchmarks. We visualize a comparison of our dataset with existing 3D and spatial benchmarks (Table 1). 3D understanding is vital for various computer vision tasks, including segmentation (jatavallabhula2023conceptfusion; hong20233d; kerbl20233d; nie2020total3dunderstanding; kundu2022panoptic; wang2023octformer), object localization (qi2019deep; sajjadi2022object; kamath2021mdetrmodulateddetection), tracking (agia2022taskography; kurenkov2023modeling; li2022pre), fine-grained scene captioning (chen2021scan2cap; su202125dvisualrelationshipdetection), open-vocabulary classification (chen2020scanrefer; hong20233d; achlioptas2020referit3d; shao2019objects365), and question answering (ye20213d; azuma2022scanqa; linghu2024multimodalsituatedreasoning3d). Most use 3D scans of real-world data (hong20233dllm), but 3d scans are expensive to recompute for dynamic scenes and most MLMs are trained to input 2D images-hence we focus on 3D spatial reasoning on 2D inputs. In light of potential applications in embodied AI (duan2022survey), we align our dataset to be compatible with physics simulators, allowing us to go beyond datasets which only leverage image annotations (caesar2020nuscenesmultimodaldatasetautonomous) and have static spatial reasoning (fu2024blink; duan2022pip; chen2024spatialvlm; kamath2023whatsupvisionlanguagemodels; su202125dvisualrelationshipdetection; liu2023visualspatialreasoning; goyal2020rel3d; yang2019spatialsenseadversariallycrowdsourcedbenchmark; du2024embspatialbenchbenchmarkingspatialunderstanding).
Synthetic training data. Many have investigated whether synthetic information can boost reasoning in real environments geng2024unmet, finding that synthetic images sometimes help in classification (chen2021contrastivesyntorealgeneralization), semantic understanding (mishra2022task2sim), correcting biases (qraitem2023fake), and teaching navigation capabilities to embodied AI (ehsani2024spoc; silwal2024we). Inspired by this, our work explores if synthetic data with perfect 2D/3D information can improve spatial reasoning.
Vision-language models. Our task is heavily influenced by the emergence of multimodal foundation models (radford2021learning; jia2021scaling; wang2022omnivl; yuan2021florence; fu2021violet; wang2022internvideo; ye2023mplug). VLMs have been adapted for a wide range of downstream image (tsimpoukelli2021multimodal; driess2023palm; zhu2023minigpt; li2022pre; alayrac2022flamingo; dai2024instructblip) and video (zhang2023video; maaz2023video; tan2024koala; wang2024internvideo2; li2023videochat; luo2023valley) understanding tasks. Many tasks require spatial understanding of the scene (brohan2023rt; zhang2024feedback). Multiple recent works (fu2024scene; hong20233d; cho2024language) have noted their weakness in 3D pose and location estimation. Adjacent to works like ray2023cola; zhang2024task; hsieh2023sugarcrepe; thrush_winoground_2022 that point to deficiencies in understanding compositions of spatial relationships with objects and attributes, we explore if perfect 3D information in synthetic images can improve spatial understanding.
3 SAT: Spatial Aptitude Training
Our goal is to improve the spatial reasoning capabilities of MLMs. Since obtaining varied 3D annotations with interactive movements of the camera and objects on real images is expensive and tedious, we procedurally generate photo-realistic environments and automatically curate training data. The resulting data generation pipeline, SAT, serves as both instruction-tuning data for MLMs and as a benchmark to test the dynamic spatial reasoning capabilities not present in existing benchmarks.
In total, our training dataset contains 175175K questions across 2222K procedurally generated scenes from ProcTHOR-10K dataset of indoor apartment buildings. Next, we outline the generation process for the two kinds of spatial questions—static and dynamic.
Figure 2: Method of generating our SAT dynamic data: we take actions in a 3D simulator and check the 3D locations of assets. We generate template-based QA pairs based on how the 3D nature of the scene changes with actions taken.
3.1 Generating Static Spatial QAs.
Aligning with contemporary benchmarks for spatial reasoning, we first generate instruction-tuning data for static spatial reasoning that deal with relative relations of objects. Overall, we generate 127127K static spatial QA pairs across 88K images across the following types:
Relative spatial relations. We generate questions about the relative location of one object in the scene to other objects. We form two kinds of questions—(1) judging if object X is to the left, right, above, or below object Y. For example, Is the wine bottle to the left or right of the plate? (2) judging if object A or B is closer to another object C. For example, Which object is closer to the wine bottle—the cup, or the plate? Given the camera parameters, we first project the objects’ poses into the camera coordinate system and then generate the corresponding answers. More details about the camera coordinate normalization are in the appendix.
Relative Depth. We generate questions about judging whether object X is closer to the camera than object Y by calculating the distances to the objects in the simulator. For example, Is the wine bottle closer to the camera than the plate?
Count. Since many MLMs (goyal2017makingvvqamatter) struggle with counting, we include counting questions using the number of object instances from the metadata of our simulator.
3.2 Generating Dynamic Spatial QA
Grounded in spatial cognitive tests (anderson2013role; brucato2023measuring; vasilyeva2012development), we outline five different complex tasks that require reasoning about egocentric movements, object movements, and allocentric perspectives. We generate 4848K QAs on 1313K images. The high-level idea behind generating such QAs is illustrated in Figure 2. Given a frame in a simulated environment, we take an action and formulate QAs based on how the 3D orientation of objects changes based on the action taken. Below, we outline the specific approach for each of the question types.
Egocentric Movement. This is based on the “moving room test” anderson2013role, a fundamental test designed to assess and improve spatial cognitive development in children. This test aims to measure if an agent can judge how they moved given two frames. This is useful beyond just measuring spatial cognition since this task can help judge navigation trajectories from egocentric videos in embodied applications. We take a random action from the choices of rotating left or right by a random angle, or moving left/right by a random distance. Note that moving and rotating are different since a camera can be moving left while rotating right. Based on the action taken, we formulate a question of the type: How did the camera taking the video likely rotate/move? with the answer being the action sequence taken. We have 6.96.9K training image-QA pairs of this type. This is denoted as EgoM in the tables. More details of the exact simulator actions taken to generate this are in the appendix.
Object Movement. Here, we randomly choose an object and move it by a randomly chosen distance and direction, ensuring that the object is still in the frame of view. Next, we compare the updated 3D position of the object with the original position normalized by the camera coordinates to decide if the object got closer, further, more left, or more right from the original position, or if it did not move. Based on that, we form QA pairs of the type: Did any of the objects move from the first frame to the second frame? with the answer being the way the object moved. Note that sometimes objects may move in conjunction with camera movement. To answer the questions accurately, the agent needs to learn to distinguish between egocentric movement and objects moving. We have 6.96.9K training image-QA pairs of this type. This is denoted as ObjM in the tables.
Allocentric Perspective. Inspired by a spatial cognitive test for humans and animals brucato2023measuring, this test checks if the agent is able to take the perspective of another viewer and judge the relative locations of objects according to the other viewer. To make such reasoning QAs, we first choose a 2D point in the scene and mark it as “X". Next, we teleport that agent to the 3D location corresponding to “X" (determined by ray tracing). We check the relative positions of objects according to the camera view from “X” (similarly as described in 3.1. We make QA’s of the type: For someone at the mark ‘X’ facing left/right by 90 degrees, would the <object> be to their left or right? We have 66K training image-QA pairs of this type. This is denoted as Pers in the tables.
Goal Aiming. Aiming is a prerequisite for efficient navigation to objects franz2000biomimetic, a fundamental spatial cognitive capability. Hence, we design QAs that check how well agents can aim to the desired object. We pick a random object and calculate the angle of the object to the camera using the 3D location of the object and camera, assuming looking forward is 0 degrees (exact equations in the supplementary). Based on the angle, we formulate questions of the type: I need to go to the countertop. Which direction should I turn to face it? Since precise angles are hard to judge from a single image, we formulate answers as choices of rough angles to turn towards the left or right. We have 6.86.8K training image-QA pairs of this type. This is denoted as Aim in the tables.
Action Consequence. Here, the agent needs to reason about how the spatial relationships change when it takes a movement action, inspired by how humans can reason about the consequences of their movements in an environment franz2000biomimetic. Here, we show the first frame and ask the agent to judge if we would move closer/further, or look towards or away from an object if it took that action. e.g., If I rotate left by 90 degrees and move forward, would I move further from the sofa? Note that in most cases, moving forward would get us closer to an object. To make the distribution of answers even, we sometimes rephrase the question as to whether we would be facing the object or not. We have 1515K training image-QA pairs of this type. This is denoted as EgoAct in the tables.
Real-image Test Set. Existing image spatial datasets only test for static relationships and not the above-mentioned dynamic capabilities. A recent video spatial dataset (yang2024thinking) also focuses on static object relations in a video input where the scene doesn’t change. Hence, we construct a small yet challenging test set of dynamic QAs that focus on how relationships change given ego or object motions and varying perspectives. We use expert annotators to annotate 150 image-question-answer pairs ($\sim 30$ each of the above dynamic types). We also generate a larger synthetic test set—$647$ object movement, $647$ egocentric movement, $592$ goal aim, $1336$ action consequence, and $778$ perspective questions on $805$ images for analysis on individual splits.
3.3 Finetuning MLMs with SAT
We provide the MLM with the required image(s) and the question and ask it to choose between two answers—a correct answer and a distractor answer, following existing benchmark standards (fu2024blink; cai2024temporalbench). The distractor answer is generated by switching the correct answer to the opposite/distractor word (e.g., left $\rightarrow$ right, or “did not move”). To prevent catastrophic forgetting, we mix in some of the original pre-training data during tuning from the LLaVA Instruct-tune dataset (liu2024visual). More tuning and prompt details are in the appendix.
4 Experiments
Using our SAT data as a training set and existing spatial benchmarks along with our SAT real-image dynamic test set, we investigate what encourages spatial aptitude in MLMs.
Setup. We use two popular open-source MLMs: LLaVA-1.5-13B (liu2024visual) and LLaVA-Video-7B (zhang2024video) for our experiments. We report the standard accuracy metric used for question-answering evaluations by checking if the predicted answer matches the GT answer. We notice better performance for off-the-shelf MLMs when we provide the options in text (e.g., Choose between right or left) as opposed to option numbers (e.g., Choose between option A or B). Hence, we sometimes report higher performances for the baselines than those reported by the original papers. Below, we present the key experimental questions we wish to ask.
Table 2: Both large proprietary and open-source MLMs struggle on our dynamic SAT spatial test set, including strong models like Gemini1.5-pro and spatially-tuned Robopoint. SAT training improves LLaVA models significantly on SAT test set.
SAT Real SAT Synthetic
Avg EgoM ObjM EgoAct GoalAim Pers Avg
a. GPT4-V 50.7 54.7 32.7 52.0 50.5 34.2 44.8
b. GPT4-o 57.5 61.5 33.2 47.6 67.5 37.5 49.4
c. Gemini1.5-flash 57.6 67.1 33.1 52.9 64.0 32.7 50.0
d. Gemini1.5-pro 64.8 57.7 29.8 55.5 56.9 49.5 49.9
e. Robopoint-13B 46.6 50.2 69.4 48.8 72.6 25.5 53.3
f. LLaVA-1.5-13B 41.6 46.6 73.8 49.7 45.6 39.9 51.1
g.   + SAT 54.9 61.7 90.2 91.4 96.8 98.5 87.7
$\Delta$ Improvement +13.3 +15.1 +16.4 +41.7 +51.2 +58.6 +36.6
h. LLaVA-Video-7B 53.5 56.4 82.7 48.0 52.9 47.1 57.4
i.   + SAT 63.4 79.6 80.4 85.3 56.4 88.4 78.0
$\Delta$ Improvement +9.9 +23.2 -2.3 +37.4 +3.5 +41.3 +20.6
Table 3: SAT tuning improves performance of open MLMs on existing static spatial relationship benchmarks—often outperforming large proprietary and spatially-tuned Robopoint.
CVBench (tong2024cambrian1fullyopenvisioncentric) BLINK (fu2024blink)
Count 2DRel 3DDep 3DDis Avg MV RelDep SpRel Avg
a. GPT4-V 62.4 71.1 79.8 68.3 70.2 55.6 59.7 72.7 62.7
b. GPT4-o 65.9 85.7 87.8 78.2 78.9 58.6 69.4 82.5 70.2
c. Gemini1.5-flash 66.0 76.9 75.3 68.3 71.4 51.1 62.9 62.9 59.0
d. Gemini1.5-pro 70.4 85.2 82.4 72.8 77.4 36.8 70.2 70.6 59.2
e. Robopoint-13B 53.6 79.4 74.7 71.3 69.1 48.1 51.6 75.5 58.4
f. LLaVA-1.5-13B 58.2 46.6 53.0 47.8 51.7 45.1 56.4 69.9 57.1
g.   + SAT 61.5 89.7 80.7 73.0 75.6 44.4 76.6 72.7 64.6
$\Delta$ Improvement +3.3 +43.1 +27.7 +25.2 +23.9 -0.7 +20.2 +2.8 +7.4
h. LLaVA-Vid-7B 59.3 77.0 71.3 54.7 65.2 39.1 55.6 75.5 56.7
i.   + SAT 66.2 81.2 88.2 79.3 78.4 48.1 66.1 73.4 62.6
$\Delta$ Improvement +6.9 +4.2 +16.9 +24.6 +13.2 +9.0 +10.5 -2.1 +5.8
Table 4: SAT also improves spatial performance on videos, VSI-Bench (yang2024thinking)
Rel Dist Rel Dir Rt. Plan App. Order MC Avg
a. GPT4-o 37.0 41.3 31.5 28.5 34.6
b. Gemini-1.5-flash 37.7 41.0 31.5 37.8 36.9
c. Gemini-1.5-pro 51.3 46.3 36.0 34.6 42.1
d. LLaVA-Video-7B 43.9 42.0 33.5 32.3 37.9
e.   + SAT 47.9 39.6 38.7 40.6 41.7
$\Delta$ Improvement +4.0 -2.4 +5.2 +8.3 +3.8
Table 5: SAT also improves out-of domain spatial relation (such as outdoor and video reasoning with humans on MME-RealWorld-Lite) despite having only indoor scenes.
Model Avg Motion Interaction Position Reasoning Perception
LLaVA-Vid-7B 35% 26.8% 24.5% 34% 36.8% 35.0%
+ SAT 39% 33.2% 31.5% 42% 39.7% 39.6%
$\Delta$ Improvement +4% +6.4% +7% +8% +2.9% +4.6%
Table 6: Performance on some traditional spatial benchmarks- VSR (liu2023visualspatialreasoning), GQA-Sp (hudson_gqa_2019)- and other VQA benchmarks- GQA (hudson_gqa_2019), OK-VQA (marino2019okvqavisualquestionanswering), VQAv2 (goyal2017makingvvqamatter)- not focused on spatial relationships. Our spatial tuning remembers pre-training commonsense on other capabilities while improving spatial reasoning.
VSR GQA-Sp GQA OK-VQA VQAv2
a. LLaVA-1.5-13B 65.9 53.1 78.6 30.7 60.5
b. + SAT 70.4 55.8 71.8 35.1 60.9
4.1 Can SAT improve spatial reasoning in MLMs?
To evaluate the effect of SAT training on spatial performance, we use 4 existing static spatial benchmarks—CVBench (tong2024cambrian1fullyopenvisioncentric), BLINK (spatial subsets) (fu2024blink), Visual Spatial Relations (VSR) (liu2023visualspatialreasoning), and GQA-Spatial (hudson2019gqanewdatasetrealworld). For dynamic reasoning, we use our SAT test sets on both real and synthetic images. Since SAT real is small, we only report the overall average and report individual split performance on our synthetic test set for a more detailed analysis. Finally, we evaluate the effect of SAT tuning on a recent long-video spatial benchmark, VSI-Bench (yang2024thinking), by tuning LLaVA-Video-7B (zhang2024video).
Open MLMs struggle on spatial reasoning. While proprietary MLMs perform better on static reasoning, they still struggle with dynamic reasoning. In Tables 3 and 2, we see that open-source MLMs (rows f and h) struggle on both static relationships and dynamic reasoning. Larger closed-source models like GPT4-o (achiam2023gpt), Gemini-1.5-pro (team2023gemini) and spatially-tuned Robopoint (yuan2024robopointvisionlanguagemodelspatial) are better at static reasoning (in Table 3 row b, d), but still struggle on dynamic QAs (Table 2 rows a-e). On SAT synthetic, aiming at the goal is generally easier for spatially stronger models like RoboPoint and GPT4-o since it also mostly requires judging object position.
Training on SAT significantly improves static spatial reasoning on real images for MLMs. As shown in Table 3, tuning on SAT improves performance on static spatial questions for both LLaVA-1.5 and LLaVA-Video models by significant amounts (rows f vs g and h vs i). Comparing rows a-e, we observe that SAT tuning makes both LLaVA models match/outperform some closed-source models on zero-shot performance on CVBench and BLINK. Compared to an existing spatially-tuned baseline, we outperform RoboPoint-13B (yuan2024robopointvisionlanguagemodelspatial) (row e). This shows promise that SAT instruction-tuning may push performance further for these stronger models.
Training on SAT improves dynamic spatial reasoning on real images. Tuning on simulated SAT improves performance on SAT real test set as shown in Table 2 (column 1) by $12\%$ for LLaVA-1.5 and by $10\%$ for LLaVA-Video. This shows strong sim-to-real transfer. Once again, our strongest model matches Gemini-1.5-pro and outperforms GPT models.
Camera movement and out-of-domain relations are hard to improve BLINK SpRel has abstract relationships (especially with people) not present in our synthetic data (e.g., “looking away”, “surrounding”). Hence, in Table 3, gains are modest on SpRel (row f vs g), with LLaVA-Video (row h vs i) showing no improvement. Camera movements are also hard (BLINK MV). While more investigation is needed, we believe this may be due to translation invariance in ViT (dosovitskiy2021image) leading to minimal feature changes between images with subtle camera movement.
Training on SAT improves spatial reasoning in videos. Even though SAT only has up to 2 frames in its training data, we evaluate its impact on longer videos on yang2024thinking. The results are shown in Table 4. We see that SAT training helps significantly on route planning capabilities (column Rt. Plan) while also improving other spatial splits. This shows promise that SAT tuning can benefit embodied AI applications.
SAT training also improves on outdoor videos including humans. SAT dataset is rendered using indoor simulators without any humans. However, when trained on SAT, we see improvements on outdoor video reasoning that includes humans on MME-Realworld-Lite (zhang2024mme) as shown in Table 5. While MME-RealWorld has numerous splits related to reasoning and perception, we highlight some of the splits where we see the most improvements when tuned on SAT along with the overall average (Avg column) and the average for all “Reasoning” and “Perception”-based splits.
Our SAT-tuned model maintains its non-spatial reasoning capabilities. We run an evaluation on some standard VQA benchmarks—namely, GQA (hudson2019gqanewdatasetrealworld), VQAv2 (goyal2017makingvvqamatter), and OK-VQA (marino2019okvqavisualquestionanswering) (9K image-QA pairs, random 3K from each). We maintain overall performance on them compared to the off-the-shelf LLaVA (liu2024visual) as shown in Table 6. This suggests that we remember pre-trained vision-language commonsense while adding stronger spatial capabilities.
SAT synthetic is a useful diagnostic test set The lack of improvement on object relations (SpRel and 2DRel) in LLaVA-Video is also reflected by a lesser gain on the GoalAim (row h vs i, Table 2) on SAT synthetic test, where the model needs to judge object position. Similar for SAT EgoM (Table 2) and BLINK MV (Table 3) dealing with camera movements-LLaVA-Video improves more than LlaVA on both. Hence, despite being synthetic, SAT synthetic is also a useful diagnostic test set.
Table 7: Table showing the effectiveness of high-quality but simulated SAT (row d) over real-image but noisier pseudo-annotated spatial data (row b, c). Dynamic spatial reasoning data (row e) further improves static spatial performance.
SATReal CVBench BLINK VSR
Count 2D Rel 3DDep 3DDist Avg MV RelDep SpRel Avg
a. LLaVA-1.5 41.6 58.2 46.6 53.0 47.8 51.7 45.1 56.4 69.9 57.1 66.0
b. + GQAPseudo 45.1 63.1 64.0 41.2 48.7 54.5 53.4 48.4 57.3 53.1 62.3
c. + VSR/25VRD 45.4 60.8 64.2 54.3 44.7 55.9 3.0 61.3 69.9 44.7 67.9
d. + SAT Static 46.0 59.5 81.7 72.5 54.2 66.4 55.6 66.9 66.4 63.0 68.0
e.  + Dynamic 54.9 61.5 89.7 80.7 73.0 75.6 44.4 76.6 72.7 64.6 70.4
4.2 What are effective types of spatial training data?
Sources of instruction tuning Since contemporary work on spatial understanding uses pseudo annotations (chen2024spatialvlm; cheng2024spatialrgptgroundedspatialreasoning), we wish to compare to such baselines as well. Since their datasets are not easily available yet, we reproduce a similar instruction-tuning set using pseudo annotations to create spatial QAs. We keep formats the same as SAT to ensure the performance differences are attributed to the source of spatial information and not the question formats.
  • GQAPseudo: We create SAT-Static like instruction-tuning data by inferring the depth (depthanything) on images from GQA (hudson2019gqanewdatasetrealworld) and VisualGenome (krishna2017visual) and the 2D bounding box annotations. We filter out potential incorrect relationships using simple heuristics (more details in the appendix). We create 225225K static spatial image-QA tuples. Generating pseudo-annotated dynamic QAs is not possible since we cannot easily infer the result of actions on real images, hence, we only generate static QAs.
  • VSR/25VRD: We use spatial relationship datasets like VSR (liu2023visualspatialreasoning) and 2.5VRD (su202125dvisualrelationshipdetection) to produce more real spatial QAs. These datasets contain relationships such as “touching”, or “behind”, allowing us to formulate QAs such as “Is the cat touching the sofa?” We combine VSR and 2.5VRD to create 200200K image-QA tuples.
(a) Improvement over LLaVA-1.5-13B
(b) Improvement over LLaVA-Video-7B
Figure 3: Simulation-only training vs mixing in some original pre-training data. Sim-only is very effective, mixing in pre-training data improves further.
Figure 4: Ablations with training data size. We see that performance generally scales with data, with dynamic showing better scaling.
SAT simulated data outperforms pseudo-annotations or human annotations. In Table 7, we see that pseudo-annotations can improve performance on certain splits like Count and 2DRel on CVBench (row b). However, noise in annotations (e.g., bounding boxes often not accounting for the entire object) leads to errors in judging depth of the entire object relative to other objects (e.g., a part of the annotated “lake” may be in front of a “tree”, but the majority of it may lie behind it). Following more complex careful curation similar to cheng2024spatialrgptgroundedspatialreasoning may improve this performance, which we leave to future work since their data is not yet easily available. Human-annotated spatial relations (using su202125dvisualrelationshipdetection; liu2023visualspatialreasoning) tend to perform well in-domain on VSR test. While the higher diversity of human-annotated relations also helps prevent forgetting on BLINK SpRel (row c vs d), they are finite and not easily composable to generate varied 3D and dynamic data. Hence, we observe no improvements on SAT dynamic and minimal gains in 3D perception splits.
Adding dynamic QAs further improves performance over just static QAs. We notice improvements on both dynamic and static performance when we add dynamic data as noted in Table 7 (row d vs e). However, BLINK MV remains challenging and we observe no performance gain. While static-only seems to perform better on BLINK MV, we observe that simply predicting “moved right" all the time achieves that performance. We also notice gains if dynamic mixed data matches the amount of static-only (in appendix Table 10).
SAT training scales with more data, especially when dynamic splits added. Since SAT is generated procedurally, we wish to explore whether spatial performance of MLMs scale with more data. Hence, we vary the amounts of training data (train for 1 epoch). As shown in Figure 4, we see that while performance tends to saturate with static-only training (likely due to overfitting to simpler relationships), the performance scales positively with dynamic reasoning questions mixed in.
Mixing in some original pretraining data when training in simulation helps performance We check whether simulation data alone suffices to impart spatial reasoning to MLMs. For sim-only training, we use our full SAT static+dynamic data. We compare this with sim+real training, where we mix in random samples of the original pretraining data $60\%$ of the time during the training along with our SAT data. For both cases, the model sees the same amount of new SAT spatial data. As shown in Figure 3, surprisingly, simulation alone is still effective at improving the spatial awareness on real images. However, mixing in pretraining data improves further.
5 Discussion
Limitations. We instruction-tune two kinds of MLMs for spatial reasoning, an image-only LLaVA (liu2024visual) and a video LLaVA-Video (zhang2024video). While we remember pretraining commonsense as noted in Table 6, we haven’t explored improving other capabilities like math and science reasoning yue2023mmmu. The scope of the paper, however, is to analyze what kinds of data improve spatial performance, and not a large-scale training of a new MLM. Additionally, further analysis on more MLMs (dubey2024llama; deitke2024molmopixmoopenweights) might be beneficial.
Future work. Although our study focuses on evaluating the spatial reasoning capabilities of MLMs, it can be extended in various avenues. For instance, to determine the kinds of embodied applications that benefit from improved complex spatial reasoning. We do see gains in accuracy in route planning in VSI-Bench (yang2024thinking), which should help embodied navigation. As a more direct evaluation, we also check the action prediction accuracy (from a choice of going left/right/forward) for a given frame on the SPOC EasyObjectNav benchmark (ehsani2024spoc). Our model (SAT Dynamic) scores an accuracy of $51\%$ compared to $40\%$ for training only on static spatial questions. This suggests that embodied navigation might benefit from improved dynamic spatial reasoning. We leave a more thorough evaluation in this direction as future work. Further, leveraging the interactive nature of our scenes with perfect 3D information could facilitate more explorations in dynamic and chain-of-thought causal reasoning. Our causal annotations in language and their corresponding image frames could also advance language-controlled world models genie3, a direction we leave for future work.
Conclusion. We propose a training and testing dataset of dynamic motion-based spatial tasks that go beyond simpler static perception of relationships in existing datasets. This improves MLMs on numerous spatial benchmarks while maintaining pre-trained commonsense. We hope that SAT paves the way for bridging the gap between passive perception and active interaction reasoning in MLMs, making them more suitable for deployment in real-life applications.
Appendix A Appendix
Acknowledgements
This project was partially funded by Amazon Science and NSF award number 2120322. We wish to thank Devdeep Ray and Jang Hyun Cho for helpful discussions on 3D and depth estimation. The contents of the paper do not represent the views of the US Government or the funding agencies.
A.1 Discussion
Limitations. We instruction-tune two kinds of MLMs for spatial reasoning, an image-only LLaVA (liu2024visual) and a video LLaVA-Video (zhang2024video). While we remember pretraining commonsense as noted in Table 6, we haven’t explored improving other capabilities like math and science reasoning yue2023mmmu. The scope of the paper, however, is to analyze what kinds of data improve spatial performance, and not a large-scale training of a new MLM. Additionally, further analysis on more MLMs (dubey2024llama; deitke2024molmopixmoopenweights) might be beneficial.
Future work. Although our study focuses on evaluating the spatial reasoning capabilities of MLMs, it can be extended in various avenues. For instance, to determine the kinds of embodied applications that benefit from improved complex spatial reasoning. We do see gains in accuracy in route planning in VSI-Bench (yang2024thinking), which should help embodied navigation. As a more direct evaluation, we also checked the action prediction accuracy (from a choice of going left/right/forward) for a given frame on the SPOC EasyObjectNav benchmark (ehsani2024spoc). Our model (SAT Dynamic) scores an accuracy of $51\%$ compared to $40\%$ for a model trained only on basic spatial questions. This suggests that embodied navigation might benefit from improved dynamic spatial reasoning. We leave a more thorough evaluation in this direction as future work. Further, leveraging the interactive nature of our scenes with perfect 3D information could facilitate more explorations in dynamic and causal reasoning.
A.2 Tuning details
We tune a widely used open-source MLM, LLaVA-1.5-13B liu2024visual for fine-tuning experiments. We LoRA-tune (hu2021lora) with rank $256$ and alpha $512$. We full fine-tune the LLaVA-Video-7B model since it is smaller. We see minimal differences in trends by training with more lora parameters or with full fine-tuning as shown in Figure 7. To prevent catastrophic forgetting, we randomly sample examples from the LLaVA Instruct Tuning dataset liu2024visual with $40\%$ probability while tuning with our synthetic QA pairs. We train all models to around 1 epoch of synthetic data. Due to memory constraints, we use a batch size of $8$ (using gradient accumulation). We set a small learning rate of $5e^{-6}$ (due to our small batch size) with cosine annealing with 1K warm-up steps and a weight decay of 0. Training requires two 48GB NVIDIA GPUs, while inference is possible with one GPU. To demonstrate effect on a video benchmark, we tune another widely used video MLM, LLaVA-Video-7b (zhang2024video). FOr LLaVA-Video-7b, the paper results are solely by training on SAT, without mixing any real orginal pretraining data. However, we also do compare the effect of mixing in real pretraining data as show it in Figure 4. We also show the sacling and
A.3 Dataset Details
Human Performance on our SAT We conduct a human study with experts to measure the quality of our synthetic test data which is automatically generated. We observe that spatial awareness demands more mental power since one has to pay more attention and reason about how the orientation of the scene changed or would change based on an action. We conduct an expert human study, where we ask anonymized graduate students to answer 200 randomly sampled questions from our test set using the interface showed in Figure 10. We see that humans are $92.8\%$ accurate on our SAT dataset. This is still a significant gap compared to the performance of best existing zero-shot MLM (around $53\%$). We note that zero-shot SAT accuracies on the synthetic test set are lower since synthetic image are often out of domain for MLMs. Hence, we most use SAT-synthetic to as diagnostic set to analyze performance gains after tuning the MLMs. We see tha gains on the synthetic test set are indicative of performance on the related task on real images. We will release the dataset on Huggingface. Our SAT real dataset was annotated by careful cross-checking by 4 graduate students. Our datasets will all be released publicly.
More details on dataset creation
We first take an apartment from ProcTHOR-10K and place the camera at a position where many objects are visible. We do this by randomly choosing 20 points to place the camera and then choosing the point with max objects visible.
Normalizing the camera coordinates In ProcTHOR deitke2022️, in the camera view, the $y$ coordinate is the height coordinate, which means the $y$ increases pointing upwards (e.g., the ceiling has a greater $y$ than the floor). Hence, from the bird’s eye view, the coordinates are $x$ and $z$. The rotation of the camera is such that it is always parallel to the $x$-$z$ plane. Hence the rotation is described as angle clockwise around the $y$-axis with the camera pointing to the positive z-axis as a 0-degree rotation.
Given a camera rotation, we normalize the view by translating to $(0,0)$ for $x$ and $z$ by subtracting the camera $x_{0}$ and $y_{0}$. Further, we rotate the $x$-$z$ plane such that the camera points to the positive $z$-axis.
For rotation, we use the formula:
$R=\begin{bmatrix}\cos(a)&-\sin(a)\\ \sin(a)&\cos(a)\end{bmatrix}$
Hence, the normalized $x^{\prime}$, $z^{\prime}$ for any object is computed using:
$\begin{bmatrix}x^{\prime}\\ z^{\prime}\end{bmatrix}=R\cdot\begin{bmatrix}x-x_{0}\\ z-y_{0}\end{bmatrix}$
The $y$ value remains unchanged since it is the height, which is not affected since we do not change the camera height.
Hence, finally, $x^{\prime}$ goes negative to the left and positive to the right, $z^{\prime}$ goes positive towards the depth and $y$ goes positive upwards from the floor level. We use the values of $x^{\prime}$ and $z^{\prime}$ to calculate relative relationships (left, right, in front of, and behind) as described below.
SAT Static Spatial QAs
Relative spatial relations. For instance, if the value of $x^{\prime}$ for “chair” is lower than that of “table”, the chair is to the left of the table. We can also compute the distance between objects. We randomly choose 3 objects. We compute the pairwise distances using their $(x,y,z)$ 3D coordinates. Based on whether object 1 is closer or further to object 2, we make QAs like "Is the couch closer to the lamp or the table?"
Relative Depth. Similarly, if the value of $z^{\prime}$ for, say, “lamp” is greater than that of “couch”, we say the “lamp” is further away from the camera than the couch.
SAT Dynamic Spatial QAs
Egocentric Movement. We first choose an image frame. Next, we first choose to rotate left or right from angles 20, 30, 40, 50, 60 chosen randomly. We use the controller.step(action=’RotateRight’, degrees=angle) function in the AI2THOR Kolve2017AI2THORAn platform. Next, we move forward with probability $0.5$ by a random distance from 20 to 40 centimeters (controller.step(action=’MoveAhead’, moveMagnitude=dist). We capture the next frame from this final position. We also formulate camera movement questions. Note that camera movement and rotation are separate since a camera can be moving left while rotating right. For camera movement, we follow a translate the camera left/right in a random direction by a random distance. Specifically, we first rotate by an angle between 45 to 90 degrees to the left/right, then move forward by a random distance and then rotate right/left (right if previously rotated left in the first step and vice versa) by an angle between 90 and 135 degrees. Sometimes, we do not move or rotate.
Object Movement. We first choose an object visible in the scene with at least a certain bounding box area to make sure the object is not too small or non-salient. Next we decide to move the object by a random distance sampled from 0.25 to 0.5 meters in a random direction if possible. We use the PlaceObjectAtPoint function in the AI2THOR Kolve2017AI2THORAn platform. Sometimes, we do not move any object. Sometimes the camera as well as an object moves. The agent needs to disentangle the two to answer questions correctly.
Allocentric Perspective. We choose a point on the 2D image that is not too close to the margsns (in between 0.20.2-$0.8$ normalized width and height of the image). We use GetCoordinateFromRaycast action in AI2Thor Kolve2017AI2THORAn to get the 3D location of the point. We try a few points since some points may not possible to navigate to and pick one possible point. We randomly turn left or right by 90 degrees. Next, we check whether an object is left/right based on the method described in Section A.3.
Goal Aiming. We compute the angle between a randomly chosen object with the camera assuming looking straight is 0 degrees. If the angle of an object is $\alpha$, we say we have to “turn left by $\alpha$ degrees” and “right by $\alpha$ degrees” otherwise. When $|\alpha-0|\leq\epsilon$, we say we have to look “roughly straight”. We define $\epsilon=10$. Note it is very hard for a human or machine to judge the exact degrees to turn from one single image, hence, we give multiple choices where the difference in choices are between left/right and hence, the machine really only has to decide between them and not the exact angles to turn. We use the following equation to calculate the angle of an object with the camera. First we normalize the object coordinates to the camera coordinates, $(x_{0},y_{0})$, and then calculate the angle, $\alpha$.
$\begin{bmatrix}x^{\prime}\\ z^{\prime}\end{bmatrix}=R\cdot\begin{bmatrix}x-x_{0}\\ z-y_{0}\end{bmatrix}$
$\alpha=\arctan\left(\frac{x^{\prime}}{z^{\prime}}\right)$
Action Consequence. Here, we just compute at the objects we got close to or further away while taking the action for the action sequence task or the perspective task. We formulate questions based on that. Note that most of the time we get close to objects in the scene and hence we random subsample such cases.
Pseudo-annotated QAs
We use DepthAnything (depthanything) to estimate depth of the scene and bounding box annotations in GQA (hudson2019gqanewdatasetrealworld) and VG (krishna2017visual) to formulate static spatial QAs. To make spatial QAs, we first choose three objects in the scene based on which we make questions as defined in Section 3.1. We use the following heuristics to first choose the three objects to reduce noise:
  • Too small objects: We do not choose objects where the bounding box area occupies less than $10\%$ of the area. We observe this removes a lot of noisy annotations and very non-salient objects.
  • Partially occluded objects: We do not choose objects if they are partially occluded by another object (based on the bounding box annotations).
Hence, we choose three clearly visible objects in the scene, estimate it’s depth and and the center point location. Based on this 2.5D information, we formulate questions about their depth and relative locations (left/right, behind or in front of).
Evaluation details
Evaluating on VSR and GQA-Sp
For VSR (liu2023visualspatialreasoning), we have captions like “truck in front of airplane.” We re-formulate them to into quetsion-answer pairs like: “Is truck in front of airplane? Answer yes or no” to keep all evaluation formats consistent. This ensures performance differences are due to the difficulties of spatial relationships in the datasets and not the question-answering format.
Evaluating on VSI-Bench
For VSI-Bench (yang2024thinking), we only use the MC split and not the numerical splits. We believe the numerical splits requiring estimation of exact distances and dimensions from RGB images alone without camera and depth information is difficult and out of scope for SAT since SAT focuses on relationships in images and dynamic scenes and not exact depth and distances.
Table 8: Datasource for LLaVA-Video on static and dynamic
SAT Real CVBench BLINK
Count 2DRel 3DDep 3DDist Avg MV RelDep SpRel Avg
LLaVA-Video 53.5 59.3 77.0 71.3 54.7 65.2 39.1 55.6 75.5 56.7
+ SAT Static 51.6 66.2 85.7 90.5 84.3 81.2 54.9 62.9 73.4 63.7
+ SAT Dynamic 63.4 66.2 81.2 88.2 79.3 78.4 48.1 66.1 73.4 62.6
Table 9: Datasource for LLaVA-VIdeo on long video VSI-Bench. Row b and c are trained purely on simulations, while row d is simulation and orignal pre-training data (LLaVA-OV data) mixed in during training to prevent forgetting.
VSI-Bench (vid)
Rel Dist Rel Dir Rt. Plan App. Order MC Avg
a. LLaVA-Video 43.9 42.0 33.5 32.4 38.0
b. + SAT Static 43.5 43.4 34.5 30.3 37.9
c. + SAT Static + Dyn 47.7 42.0 35.6 36.1 40.4
d. + SAT Dyn + LLaVA-OV 47.9 39.6 38.7 40.6 41.7
Table 10: Effect of dynamic data keeping the data size the same as static at different scales. This is to judge the effect of types of simulated spatial data. We see that adding dynamic maintains static spatial reasoning while improving dynamic and video spatial reasoning.
SAT Real CVBench BLINK VSI-Bench
CVB 2D CVB 3D Avg Avg MC Avg
LLaVA-13B 41.58 52.95 50.40 51.68 57.13 -
+Stat 127K 40.08 69.40 76.42 73.40 59.84 -
+Stat+Dyn 127K 51.26 73.37 77.33 75.99 63.26 -
LlaVA-Video-7B 53.45 67.30 63.00 65.15 56.73 37.95
+Static 50K 49.94 74.55 85.00 79.77 59.39 38.73
+Stat + Dyn 50K 59.28 74.13 82.08 78.11 62.34 40.36
LlaVA-Video-7B 53.45 67.30 63.00 65.15 56.73 37.95
+Static 100K 51.58 75.03 87.42 81.22 63.74 37.94
+Stat + Dyn 100K 59.13 73.50 85.25 79.38 61.09 38.63
Table 11: Most frequent correct and incorrect words
Rank Correct Words Incorrect Words Incorrect Words (excluding count)
1 stationery 3 4x4
2 pan 5 iced
3 books 4 wheeler
4 my 4x4 donuts
5 into 12 cookies
6 vase 6 tea
7 straight iced poles
8 magazine 10 office
9 signs 11 game
10 flowerpot 8 dogs
Question-answer format
We frame each questions as a binary choice following existing benchmark standards (fu2024blink; tong2024cambrian1fullyopenvisioncentric), and to prevent the MLM from choosing shortcuts by eliminating obvious wrong choices in multiple choices (cai2024temporalbench).
Following LLaVA liu2024visual convention, we use <image> tokens to represent images. This is the exact prefix we use.
    A chat between a curious human and an artificial intelligence assistant.
    The assistant gives helpful, detailed, and polite answers to the human’s
    questions.
Next, we add the question prompt to the prefix:
    ###Human: <im_start><image><im_end> Human: Answer in natural language.
    Is the person facing the frisbee? Choose between the following options:
    yes, or no.###Assistant:
For questions with two images, we simply have <im_start><image><image><im_end> in the image part of the prompt.
The prefix with “A chat between a …” is something we found to be very important for LLaVA performance. Hence, we append this prefix to the question both, when tuning and testing. Further, we also found performance improvements when we specify the answer choices in text like “choose between ‘left’ or ‘right”’ than asking the model to choose an answer option letter (like A or B) or number (like option 1 or 2). We randomize the answer choice order during evaluation. Since SAT Real is small, we perform circular eval, where we prompt the model with both ordering of choices and average the performance. This reduces randomness and is robust. We also note a higher variance in performance between different training seeds on BLINK due to the small size of the dataset. However, the trends remain the same. We will release all checkpoints, the training script, and the best tips and tricks in the training schedule.
Figure 5: Some qualitative results of spatial question answering on some of our SAT Real Dynamic Benchmark
Figure 6: Some qualitative results of spatial question answering comparing baseline LLaVA on real benchmarks.
Figure 7: Effect of scaling up more tuning parameters. Tuning vision encoder is critical - showing visual features are a current bottleneck for spatial reasoning. However, tuning more LLM parameters yields incremental returns.
Figure 8: Qualitative examples of our synthetic SAT dynamic data
Figure 9: Qualitative examples of our synthetic SAT static data
Figure 10: Interface to compute human accuracy for SAT.
A.4 Extra ablations and analysis.
Analyzing the semantic concepts where models fail after SAT tuning. We wish to understand the semantic concepts where SAT tuning doesn’t suffice. Hence, we we analyze the words that most often occur in the failure cases in contrast to the correct cases and show them in Table 11. Unsurprisingly, we see a high fraction of count words with higher numbers than 2 being mostly present. If we exclude the coutn words, we see a higher fraction of outdoor words. This is also unsurprising since our simulator only has indoor scenes. Using world models and more diverse simulators to improve our dataset is an exciting area of future work.
Errors in relationship judgment compounded with lack of reasoning causes MLMs to perform worse than random on dynamic tasks For certain splits such as perspective, or overall performance on SAT real, some accuracies are below random chance (random is $50\%$). In dynamic examples, the spatial position of an object often switches when a new perspective is taken—e.g., a cup might be in the left side of the image, but switches to the right side after taking a new perspective. Weaker MLMs have bad spatial perception in general. However, even for MLMs with strong static relationship performance, they often answer what is currently visible (i.e., cup is on the left) rather than what actually happens after the new perspective (i.e., cup on the right). Examples are shown in Figure 5. For example, let’s look at the performance of a model which has strong static spatial performance - Gemini-1.5-pro. If we look at the second row example, the surfer is currently on the left of the X marked point and hence the model answers left instead of reasoning about how it might switch when viewing from that perspective. Similarly, in row 3, and in row 7 where the door is currently visible on the left. While this analysis is from qualitative analysis, a more thorough analysis is left as future work.
Adding in dynamic helps LLaVA-Video on dynamic and video spatial reasoning As shown in Tables 8 and 9, adding in dynamic data helps LLaVA-Video on video spatial reasoning (Table 9) and on dynamic SAT Real in Table 8, while maintaining performance on CVBench and BLINK. While scaling works better with mixing in dynamic as shown in Figure 4 in the main paper, we also analyze the effect of adding dynamic data while maintaining the same data size as static in Table 10. We see that for both models, adding in dynamic helps significantly on SAT Real. For LLaVA1.5 it helps overall. For LLaVA-Video, it helps on video and SAT real while maintaining performance on BLINK and CVBench.
Quality of data is important. Hence, simulations are an effective source. We wish to answer if the gains trivially come from just simply learning the formats. Hence, we run a noise baseline where we flip the answers $20\%$ of the time. We see performance degrades significantly. We see a reduction of $10\%$ on CVB and $18\%$ on BLINK when training with just $20\%$ noise. This means MLMs are not trivially learning the format of questions or object/attribute -name-based biases. Rather, the spatial answer quality is important.